# MAPS: A Multilingual Benchmark for Agent Performance and Security

### v2

@@ Abstract
Agentic AI systems, which build on Large Language Models (LLMs) and interact with tools and memory, have rapidly advanced in capability and scope. Yet, since LLMs have been shown to struggle in multilingual settings, typically resulting in lower performance and reduced safety, agentic systems risk inheriting these limitations. This raises concerns about the accessibility of such systems, as users interacting in languages other than English may encounter unreliable or security-critical agent behavior. Despite growing interest in evaluating agentic AI, existing benchmarks focus exclusively on English, leaving multilingual settings unexplored. To address this gap, we propose MAPS, a multilingual benchmark suite designed to evaluate agentic AI systems across diverse languages and tasks. MAPS builds on four widely used agentic benchmarks — GAIA (real-world tasks), SWE-bench (code generation), MATH (mathematical reasoning), and the Agent Security Benchmark (security). We translate each dataset into eleven diverse languages, resulting in 805 unique tasks and 9,660 total language-specific instances - enabling a systematic analysis of the multilingual effect on AI agents’ performance and robustness. Empirically, we observe degradation in both performance and security when transitioning from English to other languages, with severity varying by task and correlating with the amount of translated input. Building on these findings, we provide actionable recommendations to guide agentic AI systems development and assessment under multilingual settings. This work establishes the first standardized evaluation framework for multilingual agentic AI, encouraging future research towards equitable, reliable, and accessible agentic AI. MAPS benchmark suite will become publicly available upon publication.  [^1] https://huggingface.co/datasets/Fujitsu-FRE/MAPS
@@
  • Reproducibility Resources – links to the full datasets and codebase, along with licensing terms (MIT) to enable open use and distribution for research and development purposes.
  • Agentic AI Benchmark Comparison – an expanded analysis of existing performance and security agentic benchmarks across dimensions such as evaluation objective, task type, answer format, and multilingual support.
@@
  • Manual Evaluation of Translation Quality – annotation guidelines, evaluation examples, and a complete breakdown of human verification scores per dataset and language.
  • Experimental Settings – full specification of the technical setup, agent configurations, and language adaptation process used to isolate multilingual effects across tasks.
  • Extended Results – detailed multilingual performance tables for GAIA, MATH, SWE-Bench, and ASB across individual evaluation runs.
@@
Omer Hofman1, Oren Rachmil1, Shamik Bose1, Vikas Pahuja1, Jonathan Brokman1, Toshiya Shimizu2, Trisha Starostina3, Kelly Marchisio3, Seraphina Goldfarb-Tarrant3, Roman Vainshtein1 1Fujitsu Research of Europe  2Fujitsu Limited  3Cohere
@@ 1 Introduction
LLM-based agentic AI systems combine multi-step reasoning with external tools and memory to solve open-ended tasks such as code generation, web navigation, planning, and transactional services like booking and ordering (Acharya et al., 2025). By doing so, they extend to complex, real-world problems beyond standard LLM benchmarks. Since such real-world applications serve speakers of diverse languages, maintaining consistent reliability in every language becomes critical. However, since agentic behavior is grounded in LLMs, which often perform inconsistently across languages (Deng et al., 2023; Wang et al., 2023), agents may inherit these multilingual limitations as well, affecting their functionality and trustworthiness. This presents a barrier to equitable access, as non-English users may face degraded responses, incorrect tool actions, or unsafe behaviors—failures that can lead to actual harm in the real world, including erroneous transactions, data corruption, and security vulnerabilities (Zhang et al., 2024).
To assess emerging agentic systems, various benchmarks have been proposed to evaluate agent performance across a range of tasks (Mialon et al., 2023; Jimenez et al., 2023; Chang et al., 2024; Xu et al., 2024). However, these benchmarks remain English-only. In contrast to multilingual LLM benchmarks (Dang et al., 2024; Shi et al., 2022; Goyal et al., 2022), no equivalent exists for agentic AI—creating a blind spot in our understanding of cross-language performance, safety, and security.
@@
In this paper, we address this gap. We hypothesize that multilingual settings will reveal performance and security gaps in agentic systems that are not captured by the existing, English-only benchmarks. To investigate this, we introduce MAPS, a Multilingual Agentic AI Benchmark Suite for Performance and Security. MAPS is based on four established agentic benchmarks: GAIA (real-world tasks) (Mialon et al., 2023), SWE-bench (code generation) (Jimenez et al., 2023), MATH (mathematical reasoning) (Hendrycks et al., 2021), and the Agent Security Benchmark (security) (Zhang et al., 2024). These benchmarks are extended to eleven typologically diverse languages beyond English [^2] MAPS offers evaluation in twelve languages: English, German, Spanish, Portuguese (Brazil), Japanese, Russian, Italian, Arabic, Hebrew, Korean, Chinese and Hindi., by employing a hybrid machine and LLM-based translation approach Ki and Carpuat (2024) with extended verification and enhancements. In total, MAPS includes 805 unique tasks, each available in 12 language versions—including the original English and 11 translated variants—for a total of 9,660 instances. An overview of the benchmark layout is shown in Fig. 1.
@@
To demonstrate the use of MAPS and test our hypothesis, we selected a leading open-source agent associated with each of the four original benchmarks and applied it to the corresponding multilingual extension. We observed notable declines in both task performance and security when moving from English to other languages, with the severity of these drops varying by task type and correlating with the proportion of non-English input tokens, suggesting that multilingual performance interventions should be targeted based on input composition and task sensitivity. Beyond overall degradation, our findings reveal that multilingual inputs can amplify agentic vulnerabilities in safety-critical tasks, highlighting the need for multilingual risk assessment. These results empirically support our hypothesis and demonstrate the utility of MAPS for multilingual evaluation of agentic AI systems.
@@
  • To the best of our knowledge, we introduce the first multilingual benchmark suite for agentic AI, extending four widely used benchmarks into eleven typologically diverse languages for systematic performance and security assessment.
@@ 2.1 Agentic AI Benchmarks
With the rapid advancement of LLM-based agents, a diverse suite of benchmarks has been developed to assess their autonomy, tool use, planning, and memory integration (Yao et al., 2024; Xu et al., 2024; Yehudai et al., 2025). We organized these suites along three primary dimensions.
Evaluation objective: Performance-oriented benchmarks measure task completion, multi-step reasoning, and correct tool use (e.g., AgentBoard (Chang et al., 2024)), whereas security-focused suites probe robustness to adversarial inputs, jailbreaks, and unsafe behaviors (e.g., AgentHarm (Andriushchenko et al., 2024)). Agentic task scope: Full-agentic evaluations present only problem statements and expected outcomes, requiring end-to-end planning and execution (e.g., GAIA (Mialon et al., 2023)), while semi-agentic frameworks supply scaffolding, such as code templates or mock APIs, to isolate the LLM’s reasoning and tool-selection core (e.g., AppWorld (Trivedi et al., 2024). Design and evaluation characteristics: Most benchmarks span a limited set of domains (three to five use cases), typically including real-world information retrieval and navigation (e.g., AssistantBench (Yoran et al., 2024)), code generation (e.g., SWE-Bench (Jimenez et al., 2023)), reasoning and planning (e.g., MATH (Hendrycks et al., 2021), and security scenarios (e.g., Agent Security Benchmark (Zhang et al., 2024)). They enable objective measurement despite agents’ open-ended capabilities, by formulating tasks with definitive ground truth, allowing clear determination of success (Jimenez et al., 2023; Mialon et al., 2023). Further comparisons are provided in the Appendix.
While multilingual LLM benchmarks such as XTREME (Hu et al., 2020), FLORES (Goyal et al., 2022), and SIB-200 (Adelani et al., 2023) have enabled broad cross-lingual evaluation, they do not assess interactive decision-making, tool use, or task execution - the core elements of agentic systems. Thus, existing multilingual benchmarks fail to capture the complexities and vulnerabilities agents face in non-English settings - underscoring the need for multilingual agentic benchmarks.
@@ 2.2 LLMs Multilingual Limitations
Studies show that pre-trained LLMs struggle with non-English inputs, particularly in languages that are low-resource or typologically distant from English. Multilingual benchmarks (Hu et al., 2020; Liang et al., 2020) report accuracy drops when moving from English to languages such as Swahili or Nepali. These gaps stem from imbalanced training data, with $90\%$ English data, as well as challenges in tokenizing morphologically rich languages and limited multilingual fine-tuning data (Jha, 2024). Notably, even large models (e.g., GPT-4, Llama 405B) face a “cross-lingual knowledge barrier” (Hendrycks et al., 2020), (Chua et al., 2024).
LLMs also face security risks in multilingual contexts. Since most security efforts have been English-centric, models are more prone to generate policy-violating outputs when processing non-English prompts (Deng et al., 2023; Wang et al., 2023; Aakanksha et al., 2024). Multilingual safety alignment, have been shown to be effective but it also degrades general performance latency (Aakanksha et al., 2024).
Since agentic AI systems are based on these LLMs, we hypothesize that they inherit the same language-dependent performance and security limitations. These agents carry out real-world tasks , thus any inherited shortcomings can lead to severe consequences. To our knowledge, no systematic evaluation has examined these vulnerabilities in agentic systems. To address this gap, we introduce MAPS, a multilingual agentic benchmark suite.
@@ 3.1 Translation Pipeline
Our translation pipeline combines automated translation with human expert verification - emphasizing both semantic and structural integrity. We combine Neural Machine Translation (NMT) which excels at preserving format and structure (Koehn and Knowles, 2017; Aharoni et al., 2019), and LLM translation for broader high-level capabilities. This extends Ki and Carpuat (2024), adding automated quality checks, fallbacks and human expert verification. The overall process is illustrated in Fig. 2. Below is a short summary - further details are in the Appendix.
@@ 3.  $\hat{t}_{enh}\leftarrow\Phi_{\text{enh}}(s,\hat{t},L_{t})$
Human Expert Verification. To ensure this pipeline reliability across languages and task types, we conducted human verification on a representative subset of translations. Evaluation setup and results are in Subsection 3.3.
@@ 3.2 Dataset Selection and Composition
Dataset Selection. To enable robust multilingual evaluation of agentic capabilities, we construct MAPS benchmark suite from established agentic AI benchmarks. Those were selected based on four criteria: (1) strong adoption within the research community, including prior use in agentic evaluation; (2) clearly defined, closed-form answers to enable controlled evaluation; (3) sufficient difficulty to challenge open-source agents without saturating performance; and (4) practical solvability, ensuring that multilingual degradation can be measured. Accordingly, we selected four datasets spanning real-world reasoning, software engineering, mathematical problem solving, and security assessment.
Data Composition. The metadata below summarizes the multilingual extension, including language coverage, scale, and pre-processing.
Translated Languages. We selected the following eleven typologically and geographically diverse languages: German, Spanish, Portuguese (Brazil), Japanese, Russian, Chinese, Italian, Arabic, Hebrew, Korean, and Hindi. This selection enables the evaluation of agent performance across a wide range of scripts, linguistic structures, and regional user populations. Details on the specific dialects translated for each language are provided in the Appendix.
Dataset Handling. To preserve the integrity and utility of the original datasets, we applied only minimal and targeted modifications. Across all datasets, we appended translations without modifying or removing any original metadata (such as task type, difficulty level, tools available, etc). Domain-specific syntax—such as equations in MATH, code snippets in SWE-bench, and adversarial prompts in ASB—was preserved exactly, maintaining the original task structure and technical fidelity. For MATH and SWE-bench, which were not originally designed for agentic evaluation, we further applied selective filtering to retain only the most challenging tasks based on the task difficulty field. This follows common practice in prior work to align non-agentic datasets with agentic evaluation settings (Wu et al., 2023), ensuring meaningful agent evaluation while avoiding trivial cases.
Data Volume. To balance performance and security evaluation, our benchmark comprises $805$ tasks: $405$ from performance-oriented datasets (GAIA, SWE-bench, MATH) and $400$ from the Agent Security Benchmark. We selected $165$ tasks from GAIA (full validation set), $140$ high-difficulty tasks from MATH ($20$ per topic across $7$ topics), and $100$ hard and medium tasks from SWE-bench. The remaining $400$ tasks include all security-relevant prompts from ASB. Each task was translated into $11$ languages, and combined with the original English version, this results in a total of $9,660$ multilingual tasks across $12$ languages.
To validate the benchmark’s utility and examine multilingual effects, we applied a leading agent to each dataset. Full details and performance results are reported in Section 4.

### v3

@@ Abstract
Agentic AI systems, which build on Large Language Models (LLMs) and interact with tools and memory, have rapidly advanced in capability and scope. Yet, since LLMs have been shown to struggle in multilingual settings, typically resulting in lower performance and reduced safety, agentic systems risk inheriting these limitations. This raises concerns about the accessibility of such systems, as users interacting in languages other than English may encounter unreliable or security-critical agent behavior. Despite growing interest in evaluating agentic AI and recent initial efforts toward multilingual interaction, existing benchmarks do not yet provide a comprehensive, multi-domain, security-aware evaluation of multilingual agentic systems. To address this gap, we propose MAPS, a multilingual benchmark suite designed to evaluate agentic AI systems across diverse languages and tasks. MAPS builds on four widely used agentic benchmarks — GAIA (real-world tasks), SWE-Bench (code generation), MATH (mathematical reasoning), and the Agent Security Benchmark (security). We translate each dataset into eleven diverse languages, resulting in 805 unique tasks and 9,660 total language-specific instances - enabling a systematic analysis of the Multilingual Effect on AI agents’ performance and robustness. Empirically, we observe a degradation in both performance and security when transitioning from English to other languages, with severity varying by task and correlating with the amount of translated input. This work establishes the first standardized evaluation framework for multilingual agentic AI, encouraging future research towards equitable, reliable, and accessible agentic AI. https://huggingface.co/datasets/Fujitsu-FRE/MAPS
@@
  • Reproducibility Resources – links to the full datasets and codebase, to enable open use and distribution for research and development purposes.
  • Agentic AI Benchmark Comparison – an expanded analysis of existing benchmarks across dimensions such as evaluation objective, task type, answer format, and multilingual support.
@@
  • Manual Evaluation of Translation Quality – annotation guidelines, evaluation examples, and detailed human verification scores per dataset and language.
  • Experimental Settings – full setup specification, agent configurations, and language adaptation process used to isolate Multilingual Effects across tasks.
  • Extended Results – detailed multilingual performance tables for GAIA, MATH, SWE-Bench, and ASB across runs.
@@
Omer Hofman1thanks: Corresponding author: omer.hofman@fujitsu.com.thanks: Equal contribution., Jonathan Brokman122footnotemark: 2, Oren Rachmil122footnotemark: 2, Shamik Bose1, Vikas Pahuja1, Toshiya Shimizu2, Trisha Starostina3, Kelly Marchisio3, Seraphina Goldfarb-Tarrant3, Roman Vainshtein1 1Fujitsu Research of Europe   2Fujitsu Limited   3Cohere
@@ 1 Introduction
LLM-based agentic AI systems combine multi-step reasoning with external tools and memory to solve open-ended tasks such as code generation, web navigation, planning, and transactional services like booking and ordering (Acharya et al., 2025). By doing so, they extend to complex, real-world problems beyond standard LLMs. As these systems serve speakers of diverse languages, maintaining reliability across languages becomes critical. However, since agentic behavior is grounded in LLMs, which often perform inconsistently across languages (Deng et al., 2023; Wang et al., 2023), agents may inherit these multilingual limitations, reducing their reliability in non-English contexts. While prior multilingual LLM benchmarks have quantified cross-lingual degradation in generation quality(Dang et al., 2024; Shi et al., 2022; Goyal et al., 2022), its impact on agentic reasoning, planning, and tool use remains largely unexplored.
In contrast, existing agentic benchmarks span a range of tasks (Mialon et al., 2023; Jimenez et al., 2023; Chang et al., 2024; Xu et al., 2024), but remain English-only. Multilingual evaluations of agentic systems are thus essential for equitable access, as non-English users may face degradation. For instance, a non-English speaker using a banking system with an internal agent may encounter errors even if the system operates in English—issuing a non-English query alone can raise the risk of incorrect tool actions or unsafe behaviors, leading to real-world harm such as erroneous transactions and security vulnerabilities (Zhang et al., 2024).
@@
To address this gap, we introduce MAPS, a Multilingual Agentic AI Benchmark Suite for Performance and Security - designed to systematically evaluate how multilingual settings affect agent reliability and safety. MAPS is based on four established agentic benchmarks: GAIA (real-world tasks) (Mialon et al., 2023), SWE-Bench (code generation) (Jimenez et al., 2023), MATH (mathematical reasoning) (Hendrycks et al., 2021), and the Agent Security Benchmark (security) (Zhang et al., 2024). These benchmarks are extended to eleven diverse languages beyond English (Fig. 1). By employing a hybrid machine- and LLM-based translation approach Ki and Carpuat (2024) with native-speaker verification, MAPS comprises $805$ tasks, each available in $12$ language (English plus $11$ translations), totaling 9,6609{,}660 instances.
@@
To demonstrate the use of MAPS, we applied a leading open-source agent from each of the four original benchmarks to its corresponding multilingual extension. We observed notable declines in both task performance and security when shifting from English to other languages, with the severity of these drops varying by task type and correlating with the proportion of non-English input tokens, suggesting that multilingual performance interventions should target input composition and task sensitivity. Beyond overall degradation, we reveal that multilingual inputs also amplify agentic vulnerabilities in safety-critical tasks, highlighting the need for multilingual risk assessment.
@@
  • We introduce MAPS, a multi-domain, security-aware multilingual benchmark suite for agentic AI, extending four widely used benchmarks into eleven diverse languages for performance and security assessment.
@@ 2.1 Agentic AI Benchmarks
With the rapid advancement of LLM-based agents, a diverse suite of benchmarks has emerged to assess their autonomy, tool use, planning, and memory integration (Yao et al., 2024; Xu et al., 2024; Yehudai et al., 2025). These suites differ from standard LLM evaluations across three main dimensions: Evaluation objective: Performance-oriented benchmarks measure task completion, multi-step reasoning, and correct tool use (e.g., AgentBoard Chang et al. (2024)), whereas security-focused suites probe robustness to adversarial inputs and unsafe behaviors (e.g., AgentHarm Andriushchenko et al. (2024); Brokman et al. (2025); Betser et al. (2026)). Agentic task scope: Full-agentic evaluations present only problem statements and expected outcomes, requiring end-to-end planning and execution (e.g., GAIA Mialon et al. (2023)), while semi-agentic frameworks supply scaffolding, such as code templates or mock APIs, to isolate the LLM’s reasoning (e.g., AppWorld Trivedi et al. (2024)). Design characteristics: Most benchmarks span a limited set of domains, typically including real-world information retrieval and navigation (e.g., AssistantBench Yoran et al. (2024)), code generation (e.g., SWE-Bench Jimenez et al. (2023)), reasoning and planning (e.g., MATH Hendrycks et al. (2021), and security scenarios (e.g., Agent Security Benchmark Zhang et al. (2024)). They enable objective measurement despite agents’ open-ended capabilities, by formulating tasks with definitive ground truth, allowing clear determination of success (Jimenez et al., 2023; Mialon et al., 2023). Further comparisons appear in the Appendix. Despite their breadth and diversity, nearly all existing agentic benchmarks remain English-only, leaving multilingual behavior unexplored.
@@ 2.2 LLMs Multilingual Limitations
Studies show that pre-trained LLMs struggle with non-English inputs, especially in low-resource or typologically distant languages. For instance, multilingual benchmarks (Hu et al., 2020; Liang et al., 2020) report accuracy drops when moving from English to languages such as Swahili or Nepali. These gaps may stem from imbalanced training data, challenges in tokenizing morphologically rich languages and limited multilingual fine-tuning data (Jha, 2024). Notably, even large models (e.g., GPT-4, Llama 405B) face a “cross-lingual knowledge barrier” (Hendrycks et al., 2020; Chua et al., 2024). LLMs also face security risks in multilingual contexts. Since most security efforts have been English-centric, models are more prone to generate policy-violating outputs when processing non-English prompts (Deng et al., 2023; Wang et al., 2023; Aakanksha et al., 2024). These weaknesses directly propagate to agentic systems, which rely on LLMs for decision-making and tool execution.
2.3 Multilingual LLM Benchmarks
Multilingual LLM evaluation suites such as XTREME (Hu et al., 2020), FLORES (Goyal et al., 2022), and SIB-200 (Adelani et al., 2023) mainly evaluate cross-lingual understanding and generation, covering tasks like translation and cross-lingual retrieval, but do not capture the interactive capabilities that agent benchmarks target.
2.4 Multilingual Agentic Benchmarks
Several recent works have begun to explore multilingual settings in the context of agentic or agent-like systems. For example, X-WebAgentBench (Wang et al., 2025) evaluates multilingual web navigation agents, MASSIVE-Agents (Kulkarni et al., 2025) focuses on multilingual function calling, and WebMMU (Awal et al., 2025) studies multilingual UI understanding and code generation. These efforts reflect a growing recognition of the importance of multilingual interaction in agentic systems.
However, these benchmarks are typically restricted to a single domain or a narrow interaction paradigm (such as web navigation, function calling), and rely on tightly constrained, task-specific agentic scaffolds rather than full, reusable agent frameworks. As a result, they do not capture the diversity of agentic capabilities exercised in realistic deployments, nor do they evaluate security.
MAPS is complementary to these efforts but targets a different point in the design space. Rather than focusing on a single domain, MAPS extends four established agentic benchmarks spanning real-world reasoning, software engineering, mathematical problem solving, and adversarial security scenarios. It evaluates full agent frameworks, with their native tools, and planning loops, under realistic multilingual input conditions.
@@ 3.1 Translation Pipeline
Our translation pipeline combines automated translation with human expert verification - emphasizing both semantic and structural integrity. We combine Neural Machine Translation (NMT), which preserves format and structure (Koehn and Knowles, 2017; Aharoni et al., 2019), and LLM translation for broader contextual fluency. This extends Ki and Carpuat (2024), adding automated quality checks, fallbacks and human expert verification. The overall process is illustrated in Fig. 2. Below is a short summary - further details are in the Appendix.
@@ 3.  $\hat{t}_{enh}\leftarrow\Phi_{\text{enh}}(s,\hat{t},L_{t})$
Benchmark Domain / Task Type Agentic Capability Evaluated Evaluation Metric
GAIA (Mialon et al., 2023) Real-world assistant tasks (web, retrieval, reasoning) Long-horizon planning, autonomous reasoning, and external tool use (browser, code, and document analyzers) Exact or normalized match to reference answer
SWE-Bench Jimenez et al. (2023) Software engineering (bug fixing and validation) Iterative reasoning with tool feedback, repository navigation, and code-edit planning for patch validation Test-case success rate (fail-to-pass)
MATH (Hendrycks et al., 2021) Mathematical problem solving (algebra, geometry, calculus) Stepwise symbolic reasoning and scratchpad planning; optionally assisted by computational tools (e.g., Wolfram Alpha) Exact match to final solution
ASB (Zhang et al., 2024) Security and safety stress testing Policy adherence, safe tool invocation, and memory integrity under adversarial and injected prompts Attack success rate / refusal rate
Table 1: Summary of benchmarks included in MAPS. Together, they span diverse domains and evaluation goals, enabling comprehensive multilingual analysis of agentic performance and security.
Human Expert Verification. To ensure this pipeline’s reliability, we conducted human verification on a representative subset of translations. Evaluation setup and results are in Subsection 3.3.
@@ 3.2 Dataset Selection and Composition
Dataset Selection. To enable robust multilingual evaluation of agentic capabilities, we construct MAPS benchmark suite from established agentic AI benchmarks. These were selected based on four criteria: (1) strong research community adoption, including prior use in agentic evaluation; (2) clearly defined, closed-form answers for controlled evaluation; (3) sufficient difficulty to challenge open-source agents; and (4) practical solvability, ensuring that multilingual degradation can be measured. Accordingly, we selected four benchmarks spanning real-world reasoning (GAIA Mialon et al. (2023)), software engineering (SWE-Bench Jimenez et al. (2023)), mathematical problem solving (MATH Hendrycks et al. (2021)), and security assessment (Agent Security Benchmark, ASB Zhang et al. (2024)). Each dataset represents a distinct aspect of agentic behavior-planning, tool use, reasoning, and security. Table 1 summarizes their task types, evaluated capabilities, and metrics, with full dataset details provided in the Appendix.
Data Composition.
Translated Languages. We selected eleven typologically and geographically diverse languages: German, Spanish, Portuguese (Brazil), Japanese, Russian, Chinese, Italian, Arabic, Hebrew, Korean, and Hindi, covering a broad range of scripts, linguistic structures, and regional user populations. Details on the specific dialects translated for each language are provided in the Appendix.
Data Handling. To preserve the integrity of the original benchmarks, we applied only minimal targeted modifications. Translations were appended without altering original metadata (e.g., task type, difficulty level, or tool availability) and domain-specific syntax, such as equations in MATH, code in SWE-Bench, and adversarial prompts in ASB, was preserved exactly. For MATH and SWE-Bench, which were not originally agentic, we retained only the most challenging tasks by difficulty level. This follows common practice in prior work for aligning non-agentic datasets to agentic evaluation (Wu et al., 2023), ensuring meaningful agent evaluation.
Data Volume. To balance performance and security evaluation, MAPS comprises $805$ tasks: $405$ from performance-oriented datasets (GAIA, SWE-Bench, MATH) and $400$ from the ASB. Specifically, we include $165$ GAIA tasks (full validation set), $140$ high-difficulty MATH problems ($20$ per topic across $7$ topics), and $100$ medium-to-hard SWE-Bench tasks, alongside all $400$ ASB prompts. Each task was translated into $11$ languages and combined with the original English version, yielding 9,6609{,}660 multilingual instances across $12$ languages.
To validate the benchmark’s utility and examine the Multilingual Effects, we applied a leading agent to each dataset (see Section 4 for details).
@@
Design Rationale: Fixed Environments with Multilingual Inputs. MAPS models realistic deployment settings where users interact in non-English languages while the underlying environment (e.g., documentation, code, or tools) remains primarily English. Translating only the instruction side ensures: (i) System compatibility: alignment with current deployments interfacing with English-centric tools; (ii) Cross-language comparability: all languages face identical evidence and gold answers, allowing performance and security gaps to be attributed purely to language understanding rather than environmental differences. Agents may use English internally, avoiding artificial constraints. Further discussion appears in Section 7.